OpenStudy (anonymous):

tan^-1(1) = ?. show in fractional form. show and explain steps.

OpenStudy (anonymous):

Start with the definition of tangent. tan(x) = sin(x)/cos(x) or Tan = opp/adj

OpenStudy (anonymous):

answer is pi/4, but why? how do i get that? to be clear, my question is for arctan(1) = ?

OpenStudy (anonymous):

Correct. So the ratio of sin(x) to cos(x) must be 1. This means that sin(x) = cos(x). So at what angle are cos(x) and sin(x) equal?

OpenStudy (anonymous):

Also note that it must equal (+)1. So the signs of cos(x) and sin(x) must be equal. Make note of what quadrants that not only the value of sin(x) and cos(x) are equal, but where the signs are the same also.